OpenStudy (anonymous):

How do you simplify 5x^-3y^0 with no zero, negative, or factional exponents?

OpenStudy (anonymous):

is this what you want to reduce: \[5x ^{-3y ^{0}}\]

OpenStudy (anonymous):

The y is not part of the power

OpenStudy (anonymous):

\[5x ^{-3}y ^{0}\] ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

y^0=1, 5x^-3 =5/(x^3) so we have: \[5/x ^{3}\]

OpenStudy (anonymous):

any number to the 0 power is 1, and a negative exponent is the same as one divided by the positive exponent. x^-a=1/(x^a)

OpenStudy (anonymous):

Thanks dude + medal.